Since Earth is spinning, why do we land in the same place when we jump or fall?
This question has ans on many sites but i feel confused by those ans because they say 2 things
That we move along with earth in air too
But about air resistance and how air moves along earth
So pls explain the above quest in a much easy manner with considering air resistance
1
Expert's answer
2019-06-24T15:48:12-0400

We are talking about moving up or down vertically. There is no strictly vertical movement in nature. The human movement lasts for a short period of time (a few seconds). Earth rotates at 360° for 24 hours.

For 1 second the Earth rotates on


"\frac { 360° 1 s} {24 3600 s}=\frac { 1° } {240} (1)"

At height h the person for 1 s moves in a horizontal direction on distance l


"l = \frac { 2 \pi h } {360× 240} (2)"

We put h=1 m in (2):

l=7.27×10-5 m


It's a fairly small amount.